### v1

# The NTU Partitioned Matching Game for International Kidney Exchange Programs

Abstract
Motivated by the real-world problem of international kidney exchange (IKEP), [Biró et al., Generalized Matching Games for International Kidney Exchange, 2019] introduced a generalized transferable utility matching game featuring a partition of the vertex set of a graph into players, and analyzed its complexity. We explore the non-transferable utility (NTU) variant of the game, where the utility of players is given by the number of their matched vertices. The NTU version is arguably a more natural model of the international kidney exchange program, as the utility of a participating country mostly depends on how many of its patients receive a kidney, which is non-transferable by nature. We study the core of this game, which suitably captures the notion of stability of an IKEP, as it precludes incentives to deviate from the proposed solution for any possible coalition of the players.
We prove computational complexity results about the weak and strong cores under various assumptions on the players. In particular, we show that if every player has two vertices, which can be considered as an NTU matching game with couples, then the weak core is always non-empty, and the existence of a strong core solution can be decided in polynomial time. In contrast, it is NP-hard to decide whether the strong core is empty when each player has three vertices. We also show that if the number of players is constant, then the non-emptiness of the weak and strong cores is polynomial-time decidable.
1 Introduction
In kidney exchange programs (KEP), patients may swap their incompatible donors, leading to cycles of kidney transplants [5]. In order to increase the availability of compatible donors, it is desirable for countries to merge their national patient-donor pools, leading to international KEPs (IKEPs) [7, 16]. A joint pool of multiple countries helps more patients receive kidneys and can also improve the average compatibility of transplants. However, to keep the cooperation mutually beneficial for all participating countries, it is important to guarantee some form of stability and fairness, at least in the long run. Hence, it is natural to study IKEPs in the context of game theory, which provides several theoretical tools for the analysis of various forms of stability and fairness.
In the framework of cooperative game theory, an IKEP is considered as a game where the players are the countries, and each country’s main objective is to match as many of its patients to some compatible donors as possible (in a more general model, the utility of transplantations can be characterized by edge weights, but in this paper we only consider the unweighted version). Due to difficulties and dangers involved in organizing large cycles of transplantations, it is usual to allow pairwise exchanges only. Possible pairwise exchanges are modeled by a graph $G=(V,E)$, where the vertices represent the patient-donor pairs, and an edge $uv$ represents a mutually compatible exchange. The aim of the IKEP is then to determine a matching in this graph, and each country’s main goal is for this matching to cover as many of its patients (i.e. the vertices corresponding to the patient-donor pairs of the country) as possible.
Previous literature [11, 12, 14] suggested to maintain a level of fairness using a scoring system in a multi-round mechanism. In [12], the concept of fair allocations in transferable utility (TU) games is used in each round of the IKEP to get a target allocation. This approach uses the TU version of the cooperative partitioned matching game. A comprehensive analysis of the computational aspects can be found in [3].
In the present paper, we study the computational complexity of the non-transferable utility (NTU) version of partitioned matching games. We argue that the NTU version models IKEPs more naturally, as the utilities of the countries in these programs are best described by the number of their patients that receive a kidney, which is non-transferable by nature.
In the NTU setting, we are interested in solutions where no coalition of countries has an incentive to deviate from the proposed solution in order to match more of their patients. In particular, we consider the well-known concepts of strong core and weak core in NTU games. That is, we are looking for solutions where no coalition of the players (i.e. the countries) could create a solution among themselves where each of them improves (in case of the weak core), or some of them improves and none of them has fewer patients matched (in case of the strong core).
Players in a partitioned matching game may represent entities other than countries. For example, in the context of national kidney exchange programs (KEPs), hospitals can be considered as the players of the game, which gives rise to instances where each player (hospital) has only a few vertices of the graph (patients). Thus, it is meaningful to examine the computational complexity of the problem when each player has a bounded number of vertices.
In particular, it is interesting to consider the special case where each player has only two vertices, i.e., the players correspond to couples. This is a new kind of matching problem with couples, which can be a useful modeling tool outside the realm of kidney exchange. For example, take a situation where some couples get together to play a two-person sport like squash and have to form pairs that play against each other. Each person has a list of possible opponents that they are willing to play against (which may or may not include their partner). If the graph of feasible pairs has a perfect matching, then everyone can play. Otherwise, the best option for a couple is when both can play, and the worst option is when neither of them can. We want to find a pairing where no subset of couples can organize the matches better among themselves. This gives rise to an NTU partitioned matching game with partition classes of size 2. We call this special case the NTU matching game with couples. As another motivating example, we may consider a summer research exchange program where couples are student-supervisor pairs, and we want to match each student to another supervisor whose topic is interesting to the student. The best option for an original student-supervisor pair is when both are matched in the program, and the worst is when neither are matched. Note that the graph describing the possible matches is bipartite in this instance.
The different contexts mentioned above justify the study of the computational complexity both for instances where the size of each partition class is bounded, and for instances where the number of partition classes (i.e., players) is bounded. Our main results are as follows:
  • For the NTU matching game with couples, we show that the weak core is always non-empty and a weak core solution can be found in polynomial time. As the technically most involved result of the paper, we construct a polynomial-time algorithm for deciding if the strong core is empty, and for finding a strong core solution, if it exists. The proof uses a complicated analysis of the structure of paths and cycles that are alternating between edges of the graph and “edges” corresponding to the couples. An interesting consequence of our result is that the vertex sets of matchings in the strong core can be described as the common bases of two matroids (it is well known that vertex sets of maximum matchings are the bases of the so-called matching matroid; matchings in the strong core are additionally bases of a generalized partition matroid defined by the structure of alternating paths and cycles).
  • We show that if the number of players is bounded by a constant, then the problems of checking membership in the weak and strong cores, and of finding elements in the weak and strong cores are all polynomial-time solvable. These results follow fairly easily using tools from matroid theory; here we use common bases of a matching matroid and a generalized partition matroid in a different context.
  • In contrast to the above results, we show that problems where the size of each partition class is bounded by a constant are hard even for fairly small constants. In particular, deciding whether a matching is in the weak or strong core is coNP-complete even when each partition class has size at most 3.
Our results provide useful new tools for designing mechanisms for IKEPs. In particular, as most IKEPs consist of only a few countries, our results imply that we can efficiently find strong and weak core solutions for the NTU version of the partitioned matching game in those cases. Such solutions, if they exist, give desirable outcomes for exchange programs, as strategic incentives to deviate from the program can be avoided. Also, as membership of a matching in the strong or weak core depends only on the number of vertices it covers from each player, a nonempty core always contains a maximum size matching (this follows from the fact that any set of vertices coverable by a matching is also coverable by a maximum size matching). Therefore, our concepts are compatible with the most important restriction used in KEPs, namely that the number of transplants should be maximized.
In addition to the applicability of our results to kidney exchange problems, our algorithm for the problem with couples offers a new class of non-bipartite matching games where stable solutions can be efficiently found even in the presence of couples having joint preferences.
Structure of the paper
The rest of this section describes related work and introduces the necessary definitions and notation. In subsequent sections, we study the computational complexity of membership and non-emptiness of the weak and strong cores, under various assumptions on the players.
In Section 2, we show how membership in the weak and strong cores in the NTU matching game with couples can be characterized using alternating paths. We then prove that the weak core is always non-empty in this case. The strong core can be empty, but we are able to give a structural characterization and a polynomial-time algorithm for finding a strong core solution if it exists.
In Section 3, we show that all problems studied are polynomial-time solvable if the number of players is constant, even if the sizes of the partition classes can be arbitrarily large. Section 4 presents hardess results for the case where the partition classes have bounded size. We show that all problems are hard for some small constant bound, and for most problems we can prove coNP-hardness even for partition classes of size at most 3.
Related work
In [1, 2], Ashlagi and Roth considered individual rationality in the multi-hospital kidney exchange problem. In terms of NTU games, individual rationality means that there is no single blocking player. They also studied the problem of incentive-compatibility, i.e., whether hospitals are motivated to underreport their patient-donor pairs. Carvalho et al. [8] studied the properties of Nash-equilibria in the two-hospital case. Gourvès, Monnot and Pascual [10] considered a weighted bipartite graph model, where the profit of an edge in the matching is divided between the two endpoints according to a fixed ratio. They proved that it is NP-hard to decide if there is an individually rational maximum weight matching.
As mentioned in the introduction, credit-based multi-round mechanisms for IKEPs were studied in [11, 12, 14]. Solution concepts of cooperative game theory were used in [12] to find target allocations in each round. The mechanism then computes a maximum size matching that approximates this target allocation for each country. At the end of the round, the scores are updated according to whether the number of matched patients of a country is above or below the target. Computational experiments on large-scale instances were performed in [4], while [3] studied the theoretical computational complexity of the problems involved. In particular, they examined problems similar to ours in the transferable utility setting, and showed that checking core non-emptiness in TU partitioned matching games is polynomial-time solvable for partition classes of size at most 2, but co-NP-hard for partition classes of size at most 3.
We note that a previous version of the present paper, including only a subset of the results, appeared as a technical report [13]. Independently of our work, the Master’s thesis of Collette [9] also studied the NTU version of the partitioned matching game, in particular the weak core. It offered a cutting plane approach to compute a weak core solution, as well as a comparison with the rejection-proof kidney exchange mechanism of Blom et al. [6], and computational experiments.
Preliminaries
For a positive integer $k$, we use the notation $[k]=\{1,2,\dots,k\}$. The NTU partitioned matching game is defined by a graph $G=(V;E)$ and a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$ of $V$, where $m$ is the number of players. Each class of this partition corresponds to a player of the game; we say that the vertices in $V_{i}$ belong to player $i\in[m]$ (sometimes we identify the player and the set and say that player $i$ has size $|V_{i}|$). If each $V_{i}$ has size 2, then the problem is called the NTU matching game with couples.
Given a matching $M$ in $G$, the utility of $M$ for player $i$ is defined by $u_{i}(M)=|V(M)\cap V_{i}|$, where $V(M)$ denotes the vertex set of $M$, i.e. the utility is the number of vertices of $V_{i}$ covered by $M$.
A coalition $\{i_{1},i_{2},\dots,i_{k}\}$ of players is strongly blocking for the matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for every $j\in[k]$. Similarly, a coalition $\{i_{1},i_{2},\dots,i_{k}\}$ of players is weakly blocking for a matching $M$ if there exists a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $u_{i_{j}}(M^{\prime})\geq u_{i_{j}}(M)$ for every $j\in[k]$, and $u_{i_{j}}(M^{\prime})>u_{i_{j}}(M)$ for at least one $j\in[k]$.
In the NTU setting, a matching $M$ is in the weak core if there is no strongly blocking coalition for it, and it is in the strong core if there is no weakly blocking coalition. It is easy to see that if $M$ is in the strong core, then it is also in the weak core. Deciding whether a given matching $M$ is in the strong or weak core is a problem in coNP, since a blocking coalition, together with the matching $M^{\prime}$ that certifies the blocking, is a suitable witness.
The Gallai-Edmonds decomposition
The Gallai-Edmonds (GE) decomposition of a graph $G=(V,E)$ consists of three disjoint sets of vertices $A(G),C(G),D(G)$ with $V=A(G)\cup C(G)\cup D(G)$ and the following properties:
  • $C(G)$ is the set of even-sized components and $D(G)$ is the set of odd-sized components in the graph $G\setminus A(G)$,
  • Any maximum size matching covers all but $|D(G)|-|A(G)|$ vertices, which are all members of $D(G)$ and are in different components of $G[D(G)]$,
  • Every maximum size matching has the following structure: it induces a perfect matching in $G[C(G)]$, it induces an almost perfect matching in each odd component in $G[D(G)]$ (meaning at most one vertex is exposed) and every vertex $v\in A(G)$ is matched to some vertex in $D(G)$.
  • Every component $Q$ of $G[D(G)]$ is hypomatchable, that is, $Q-v$ has a perfect matching for every vertex $v$ of $Q$. Furthermore, there is a maximum size matching where $v$ is exposed.
Matroids
In Sections 2 and 3, we will use some basic facts from matroid theory. A matroid is a pair $(S,\mathcal{I})$ of a finite set $S$ and a nonempty family $\mathcal{I}\subseteq 2^{S}$ satisfying the axioms (i) $A\subseteq B\in\mathcal{I}$ implies $A\in\mathcal{I}$, and (ii) for any $A,B\in\mathcal{I}$ with $|A|<|B|$, there is an element $x\in B\setminus A$ with $A+x\in\mathcal{I}$. A set in $\mathcal{I}$ is called an independent set. By axiom (ii), all inclusion-wise maximal independent sets have the same size; these are called the bases, and their size is the rank of the matroid.
We consider two types of matroids. Given a partition $S_{1},\dots,S_{m}$ of a ground set $S$ and numbers $0\leq q_{i}\leq|S_{i}|$, the corresponding generalized partition matroid is defined as =$\mathcal{I}=\{X\subseteq S:|X\cap S_{i}|\leq q_{i}\ (i\in[m]\}$. The rank of this matroid is $\sum q_{i}$. We also define a class of matroids using matchings. Given a graph $G=(S,E)$, the matching matroid of $G$ is defined as =$\mathcal{I}=\{X\subseteq S:\exists\ \text{matching $M$ in $G$ such that }X\subseteq V(M)\}$. The rank of this matroid is twice the maximum size of a matching. See [15, Chapter 39] for further details. We use the fact that a maximum-size common independent set of two matroids can be found in polynomial time. This implies the following lemma that we will use later.
Lemma 1.
Given a graph $G=(V;E)$, a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$ of $V$, and a vector $q\in\mathbb{Z}_{+}^{m}$, it can be decided in polynomial time if there is a matching $M$ such that $|V(M)\cap V_{i}|\geq q_{i}$ for every $i\in[m]$.
Proof.
Let $\mathcal{M}_{1}$ be the generalized partition matroid given by a partition $V_{1},V_{2},\dots,V_{m}$ and numbers $q_{1},\dots,q_{m}$, and let $\mathcal{M}_{2}$ be the matching matroid of $G$. There is a common independent set of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ of size $\sum_{i=1}^{m}q_{i}$ if and only if there is a matching $M$ with $|V(M)\cap V_{i}|\geq q_{i}$ for every $i\in[m]$. ∎
2 The NTU matching game with couples
In this section, we study the case where every partition class has size at most 2. We show that in this case all four studied problems become efficiently solvable. However, deciding the emptiness of the strong core still requires a complicated approach relying on many structural results.
First of all, we may assume without loss of generality that every partition class has size exactly two; otherwise we add an isolated vertex and make it belong to that player. With this assumption, the partition classes can be considered as edges, which we call player edges. The perfect matching defined by the player edges is denoted by $M_{0}$. Sometimes we identify a player and the corresponding player edge, so the edge of player $A$ is also denoted by $A$. These edges are not part of the original graph $G=(V,E)$, although $E$ can also contain edges between the two vertices of a player.
We say that a path or cycle in $E\cup M_{0}$ is $M_{0}$-alternating if every other edge is in $M_{0}$. Note that if $uv$ is both in $E$ and in $M_{0}$, then these two edges form an $M_{0}$-alternating cycle. Given an $M_{0}$-alternating path or cycle $P$, we can consider the coalition formed by $P\cap M_{0}$, and the matching on the vertices of the coalition given by $P\cap E$. We say that $P$ is a (weakly/strongly) blocking path or cycle for a given matching $M$ if $P\cap M_{0}$ is a weakly/strongly blocking coalition for $M$ as witnessed by the matching $P\cap E$.
2.1 Weak core
As we shall see, the problems concerning the weak core are considerably simpler than those concerning the strong core. We begin with a characterization result.
Lemma 2.
A matching $M$ is in the weak core if and only if
  1. a)
    there is no $M_{0}$-alternating cycle such that every player edge in it has at most one vertex covered by $M$,
  2. b)
    there is no $M_{0}$-alternating path such that the player edges at the two ends of the path have 0 vertices covered by $M$, and the player edges in the middle have one vertex covered by $M$.
Proof.
If $M$ is in the weak core, then conditions a) and b) hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose a) and b) hold, but $M$ is not in the weak core. There is a coalition of players with a matching $M^{\prime}$ such that every player in the coalition is strictly better off with $M^{\prime}$, therefore those players who have both of their vertices covered by $M$ cannot belong to the coalition – we can delete these players from the graph. Take the symmetric difference of the remaining player edges with $M^{\prime}$. This consists of alternating cycles and paths. If there is an alternating cycle in the symmetric difference, then it contradicts a). If there is an alternating path, then the player edges at the two ends of the path have one vertex covered by $M^{\prime}$, therefore they have 0 vertices covered by $M$, which contradicts b). ∎
Theorem 3.
If every player has size 2, then we can decide in polynomial time if a given matching $M$ is in the weak core.
Proof.
We can check the conditions of Lemma 2 in polynomial time. We restrict the graph to the players who have at most one vertex covered by $M$.
We can check if a given player edge is in an alternating cycle, since it is equivalent to checking if, after the deletion of this edge, the remaining graph (the original graph $G$ together with the player edges) has a perfect matching. Hence, condition a) can be checked by iterating through the players.
We can also check whether there is an alternating path between two given player edges, knowing that these do not belong to an alternating cycle, since this is equivalent to checking if by deleting these two edges the remaining graph has a matching that covers all but two vertices (an almost perfect matching). We can do this for any two player edges that have 0 vertices covered by $M$. ∎
In fact, we can show that the weak core of the NTU matching game with couples is never empty.
Theorem 4.
If every player has size 2, then the weak core is not empty.
Proof.
We can construct a matching $M^{*}$ in the weak core the following way. Let $M_{0}$ be the perfect matching formed by the player edges as before. We check if there is an $M_{0}$-alternating cycle in the current graph. If there is, then let the edges of $E$ in the cycle belong to $M$, and delete the vertices of the cycle from the graph. Repeat this with the remaining graph, until there are no $M_{0}$-alternating cycles. Let $M$ denote the resulting matching, and let $M^{*}$ be a maximum size matching in $G$ that covers every vertex of $M$. We claim that $M^{*}$ is in the weak core, since the conditions of Lemma 2 are met. Condition a) clearly holds, since the construction deleted players that were covered twice by $M$ (and hence by $M^{*}$), and the remaining graph did not contain an $M_{0}$-alternating cycle.
Suppose for contradiction that condition b) does not hold, i.e. there is an $M_{0}$-alternating path $P$ such that the player edges at the two ends of the path have 0 vertices covered by $M^{*}$, and the player edges in the middle have one vertex covered by $M^{*}$. Let $N^{*}$ be the set of edges of $M^{*}$ that contain a vertex from the path $P$. We have $|N^{*}|\leq|P\cap M_{0}|-2=|P\cap E|-1$. Therefore, $(M^{*}\setminus N^{*})\cup(P\cap E)$ is a larger matching than $M^{*}$, contradicting the choice of $M^{*}$ as a maximum-size matching. ∎
2.2 Strong core
Our first result on the strong core is a characterization similar to Lemma 2.
Lemma 5.
In the NTU matching game with couples, a matching $M$ is in the strong core if and only if
  1. a)
    every player that is in an $M_{0}$-alternating cycle has both its vertices covered by $M$,
  2. b)
    there is no $M_{0}$-alternating path from a player that has none of its vertices covered by $M$ to a player that has at most one vertex covered by $M$,
  3. c)
    there is no $M_{0}$-alternating path such that at least three player edges in the path have at most one vertex covered by $M$.
Proof.
It is easy to see that if $M$ is in the strong core, then these conditions hold, since otherwise the $M_{0}$-alternating path or cycle would block $M$. Suppose that the conditions hold, but the matching $M$ is not in the strong core. Then there is a weakly blocking coalition of players, together with a matching $M^{\prime}$, such that there is one player that is better off with $M^{\prime}$ than $M$, and the others are not worse off. Take the symmetric difference of $M^{\prime}$ and $M_{0}$. This is a disjoint union of $M_{0}$-alternating cycles and paths. If the player that is better off by $M^{\prime}$ has 0 vertices covered by $M$, then it has at least one vertex covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path. In the latter case, the two player edges at the ends of the path are players who have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b). If the player that is better off by $M^{\prime}$ has 1 vertex covered by $M$, then it has two vertices covered by $M^{\prime}$, so it is either in an $M_{0}$-alternating cycle, which contradicts condition a), or it is in an $M_{0}$-alternating path (and it is not at the end of the path), but the two player edges at the ends of the path are players that have one vertex covered by $M^{\prime}$, so they have at most one vertex covered by $M$, which contradicts condition b) or c). ∎
Theorem 6.
In the the NTU matching game with couples, we can decide in polynomial time if a given matching $M$ is in the strong core.
Proof.
We need to show that we can check the conditions given by Lemma 5 in polynomial time. We can check if a given player edge is in an $M_{0}$-alternating cycle, since it is equivalent to checking if, after the deletion of this edge, the remaining graph (the original graph $G$ together with the player edges) has a perfect matching. We can do this for every player edge with at most one vertex covered by $M$.
We can check whether there is an $M_{0}$-alternating path between two given player edges, knowing that these do not belong to an $M_{0}$-alternating cycle, since this is equivalent to checking if by deleting these two edges the remaining graph has an almost perfect matching. We can do this for any two player edges such that one of them has 0 vertices covered by $M$, and the other one has at most one vertex covered by $M$.
For three given player edges, such that we know that none of them belong to an $M_{0}$-alternating cycle, we can check if there is an alternating path that contains all three of them, because this is equivalent to checking if by deleting all three of these edges, the remaining graph has an almost perfect matching. We can check this for any three player edges that have one vertex covered by $M$. ∎
2.3 Existence of a strong core matching
Our aim in this subsection is to prove that it can be decided in polynomial time if the strong core is non-empty, and if so, a matching in the strong core can be found in polynomial time. These results require a careful analysis of the structure of $M_{0}$-alternating paths, so the proof will feature several structural lemmas. We start by introducing some notation.
Let $\mathcal{SC}$ denote the set of matchings in the strong core. Let $\mathcal{K}$ denote the set of players for which there is no $M_{0}$-alternating cycle through the corresponding player edge. We denote by $\mathcal{K}^{*}$ the set of players $A$ for which $\exists M\in\mathcal{SC}$: $|V(M)\cap A|\leq 1$. Of course, we cannot determine $\mathcal{K}^{*}$ without deciding whether the strong core is empty, so we will not use $\mathcal{K}^{*}$ in the algorithm, only in the proof of correctness. By Lemma 5 point a), we have $\mathcal{K}^{*}\subseteq\mathcal{K}$.
We introduce some additional notation and definitions related to the structure of $M_{0}$-alternating paths, as listed below.
  • An $M_{0}$-alternating $\delta$-path (or $\delta$-path for short) is an odd cycle and an $M_{0}$-alternating path with a common vertex $v$, where the $M_{0}$-edge adjacent to $v$ is on the path, and the cycle is $M_{0}$-alternating apart from $v$.
  • An $(A,B;C)$-$\delta$-path for $A,B,C\in\mathcal{K}$ is an $M_{0}$-alternating $\delta$-path where $A,B$ are in the cycle and $C$ is the end of the path. The vertex of $C$ at the end of the path is called the tip. Note that the roles of $A$ and $B$ are exchangeable in the definition. See Figure 1 for an illustration.
  • The two vertices of a player $A$ are usually denoted by $a$ and $a^{\prime}$. For $A\in\mathcal{K}$ and a given $M\in\mathcal{SC}$, we usually denote the path in $M\Delta M_{0}$ containing $A$ by $P_{A}$, and the two endpoints of the path by $x_{A}$ and $y_{A}$. In some contexts, $P_{A}$ will be an arbitrary $M_{0}$-alternating path that contains $A$.
Figure 1: An $(A,B;C)$-$\delta$-path. Solid edges denote the player edges, i.e. edges of $M_{0}$, while dashed edges are in $E$.
A general proof technique is to consider a graph $G^{\prime}$ obtained from $E\cup M_{0}$ by removing a single player edge $A=aa^{\prime}\in M_{0}$, and adding some additional edges, so that $M_{0}\setminus\{A\}$ is a maximum matching in $G^{\prime}$. In such a case, $G^{\prime}$ has a Gallai-Edmonds decomposition with hypomatchable odd components $Q_{1},\dots,Q_{k+2}$, where $a\in Q_{k+1}$ and $a^{\prime}\in Q_{k+2}$. We denote the player edge entering $Q_{j}$ by $S_{j}=s_{j}s^{\prime}_{j}$ ($j\in[k]$), where $s^{\prime}_{j}\in Q_{j}$. We will frequently use the following property of alternating paths: if an $M_{0}$-alternating path starts in some $Q_{i}$ ($i\in[k+2])$, does not use $S_{i}$, and ends with an edge of $M_{0}$, then the endpoint is in some $Q_{j}$ ($j\in[k+2]$). Furthermore, if $j\neq i$, then $j\in[k]$, and the path enters $Q_{j}$ at the edge $S_{j}$. This also means that such a path cannot enter an odd component $Q_{j}$ twice. See Figure 2 for an illustration. We will also use the corollary that if $b\in Q_{j}$ for some $j\in[k+2]$ and there are $M_{0}$-alternating paths to $b$ from both $a$ and $a^{\prime}$, then $j\in[k]$.
Figure 2: An $M_{0}$-alternating path in $G^{\prime}$ from $Q_{i}$ to $Q_{j}$, indicated by thick edges. Edges in $M_{0}$ are solid, while edges in $E$ are dashed.
The following lemma enables us to obtain an alternating path from the union of two paths.
Lemma 7.
Let $P$ be an $M_{0}$-alternating path from player-edge $A\in\mathcal{K}$ to player-edge $B$, with endpoints $a$ and $b$, and let $P_{A}$ be an $M_{0}$-alternating path through $A$, with endpoints $x_{A}$ and $y_{A}$. Then there exists an $M_{0}$ alternating path $P^{\prime}\subseteq P\cup P_{A}$ from $b$ to either $x_{A}$ or to $y_{A}$ (with $M_{0}$-edges at both ends) that passes through $A$.
Proof.
We assume for contradiction that there is no such path. Define $G^{\prime}=(P\cup P_{A})\cup\{x_{A}b,y_{A}b\}\setminus\{A\}$. By this assumption, there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$ in $G^{\prime}$, so we can consider the GE decomposition for $G^{\prime}$ (otherwise, as $A\in\mathcal{K}$, it contains $x_{A}b$ or $y_{A}b$ and gives such an $M_{0}$-alternating path in $G$). We may assume that $x_{A}$ is closer to $a$ than to $a^{\prime}$ on $P_{A}$. We show that the vertices $b,x_{A},y_{A}$ are in the same $Q_{i}$ for some $i\in[k]$. Indeed, there are $M_{0}$-alternating paths to all of them from $a$ or $a^{\prime}$ ending in $M_{0}$-edges, so they are in odd components of the GE decomposition, and they must be in the same $Q_{i}$ because of the edges $x_{A}b$ and $y_{A}b$. The $a-x_{A}$ path implies that $i\neq k+2$, and the $a^{\prime}-y_{A}$ path implies that $i\neq k+1$. Thus, the edge $S_{i}$ is contained in both the $a-x_{A}$ subpath and the $a^{\prime}-y_{A}$ subpath of $P_{A}$; but these two subpaths are disjoint, a contradiction. ∎
Note that the above proof works even if $y_{a}=b$; in this case, the path obtained is from $b$ to $x_{A}$. The following is an immediate corollary of Lemma 7.
Proposition 8.
Let $A,B\in\mathcal{K}$, and let $P$ be an $M_{0}$-alternating path between $A$ and $B$. Then there is no $M\in\mathcal{SC}$ such that $|V(M)\cap A|=2$ and $|V(M)\cap B|=0$.
Proof.
Suppose that $M$ is such a matching, and let $x_{A}$ and $y_{A}$ denote the two endpoints of the path $P_{A}$ containing $A$ in $M\Delta M_{0}$. Let $a$ and $b$ be the two endpoints of $P$. By Lemma 7, there exists an $M_{0}$-alternating path $P^{\prime}\subseteq P\cup P_{A}$ from $b$ to either $x_{A}$ or to $y_{A}$ (and to one that is different from $b$), that passes through $A$. But then $P^{\prime}$ is a blocking path for $M$, a contradiction. ∎
Lemma 9.
Consider an $(A,B;C)$-$\delta$-path $H$, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap C|=2$. Then there is an $(A,B;C^{\prime})$-$\delta$-path $H^{\prime}$ with the same cycle as $H$, such that the path of $H^{\prime}$ contains $C$, and $C^{\prime}$ is one end of the path in $M\Delta M_{0}$ containing $C$ (hence $C^{\prime}\in\mathcal{K^{*}})$.
Proof.
Let $P_{C}$ be the path in $M\Delta M_{0}$ that contains $C$, and let $x_{C}$ and $y_{C}$ be its two endpoints. Define $G^{\prime}$ by adding the edge $x_{C}y_{C}$ and removing $A$. Let $C=cc^{\prime}$, where $c$ is the tip of $H$. The $(A,B;C)$-$\delta$-path $H$ contains an $a-c$ path and an $a^{\prime}-c$ path, so $c\in Q_{i}$ for some $i\in[k]$. Let $P_{1}$ (respectively $P_{2}$) denote the $a-s_{i}$ part of the $a-c$ path ($a^{\prime}-s_{i}$ part of the $a^{\prime}-c$ path, respectively). The part from $s_{i}$ to $c$ is common; let this be denoted by $P$. We can use Lemma 7 for $P$ and $P_{C}$, to obtain a path $P^{\prime}\subseteq P\cup P_{C}$ from $s_{i}$ to $x_{C}$ or $y_{C}$ (we may assume $x_{C}$), that starts with $S_{i}$ and contains $C$. $P^{\prime}$ enters $Q_{i}$ only once (at its start), so it cannot intersect $P_{1}$ or $P_{2}$ before reaching $C$. It also cannot intersect them after reaching $C$ either, because that would create an $M_{0}$-alternating cycle through $C$. Let $C^{\prime}$ be the player that contains $x_{C}$. The union of $P_{1},P_{2}$, and $P^{\prime}$ gives the desired structure. ∎
It follows from the lemma that if there is an $(A,B;C)$-$\delta$-path and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$, then $|V(M)\cap B|=2$. Indeed, otherwise the $(A,B;C^{\prime})$-$\delta$-path given by the lemma would contain a blocking path for $M$ by Lemma 5.
Lemma 10.
Let $P$ be an $M_{0}$-alternating $A-B-C$ path for some $A\in\mathcal{K}$, $B\in\mathcal{K}^{*}$, $C\in\mathcal{K}$. Then there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path.
Proof.
Let $a,a^{\prime},b,b^{\prime},c,c^{\prime}$ be the vertices of $A,B,C$ in the order in which they appear in $P$. Let $M\in\mathcal{SC}$ be a matching such that $|V(M)\cap B|=1$, and let $P_{A},P_{B},P_{C}$ be the paths in $M\Delta M_{0}$ that contain $A$, $B$ and $C$, respectively; note that these paths are not necessarily distinct. We assume that one endpoint of $P_{B}$ is $b$; let the other endpoint be denoted by $x_{B}$. Let $X\subseteq V$ denote the set of endpoints of $P_{A}$, $P_{B}$, and $P_{C}$, except for $b$.
We define $G^{\prime}$ by adding a clique on $X$ and removing $B$. If there is an $M_{0}$-alternating path in $G^{\prime}$ between $b$ and $b^{\prime}$, then it must contain a new edge as $B\in\mathcal{K}$ and as we have a clique on $X$, we can assume that it contains exactly one new edge. However, deleting this edge and adding back $B$ we get an $M_{0}$-alternating path in $G$ between two players with one vertex covered in $M$ that covers both vertices of $B$, contradicting $M\in\mathcal{SC}$ by Lemma 5. Since there is no alternating path between $b$ and $b^{\prime}$ in $G^{\prime}$, we can consider the GE-decomposition. The existence of $P_{B}$ implies that $x_{B}$ is in $Q_{i}$ for some $i\in[k+1]$, since $P_{B}$ cannot enter $Q_{k+2}$. As $P$ contains a $b-a$ subpath and a $b^{\prime}-c^{\prime}$ subpath that are disjoint, it is not possible for both $a$ and $c^{\prime}$ to be in $Q_{i}$. Indeed, if $i=k+1$, then $c^{\prime}\notin Q_{i}$, while if $i\in[k]$, then $S_{i}$ cannot be on both paths (but it is included in any $M_{0}$-alternating path from $\{b,b^{\prime}\}$ that enters $Q_{i}$).
Case 1: $a\in Q_{j}$ for some $j\neq i$, $c^{\prime}\in Q_{i}$.
In this case, $i\in[k]$ and $j\in[k+1]$. Because of the clique on $X$, one endpoint of $P_{A}$ must be in $Q_{i}$. If both endpoints are in $Q_{i}$, then there is a path to $Q_{i}$ that leaves $Q_{j}$ not on $S_{j}$, so it must enter $Q_{i}$ using $S_{i}$. If one endpoint is $b$, then the path from $b$ to $a$ must contain $S_{j}$ if it exists, so we get $S_{i}\in P_{A}$ again. As $S_{i}$ is on $P_{A}$ and is also on $P_{B}$, we have $P_{A}=P_{B}$.
Let $P_{1}$ be the part of $P$ from $b^{\prime}$ to $c^{\prime}$. Let $P_{2}$ be the part of $P$ from $b$ to $s_{j}$ if $j\in[k]$, and empty if $j=k+1$. Let $P_{3}$ be the part of $P_{B}$ from $s_{j}$ (or from $b$ if $j=k+1$) to the first intersection with $P_{1}$. Such an intersection exists because $S_{i}\in P_{1}\cap P_{B}$. Note also that $P_{3}$ traverses $A$ (as $P_{B}$ cannot enter $Q_{j}$ again and $A\in Q_{j}$) and it can intersect $P_{2}$ only in $s_{j}$, because another intersection would give an $M_{0}$-alternating cycle through $A$. Thus, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is an $(A,B;C)$-$\delta$-path.
Case 2: $c^{\prime}\in Q_{j}$ for some $j\neq i$, $a\in Q_{i}$.
In this case, $i\in[k+1]$ and $j\in[k]\cup\{k+2\}$. Because of the clique on $X$, one endpoint of $P_{C}$ must be in $Q_{i}$. If both endpoints are in $Q_{i}$, then there is a path to $Q_{i}$ that leaves $Q_{j}$ not on $S_{j}$, so it must enter $Q_{i}$ using $S_{i}$, implying $P_{B}=P_{C}$. If one endpoint is $b$, then $P_{B}=P_{C}$ is obvious. We also get $j\neq k+2$.
We define $P_{1}$ as the part of $P$ from $b$ to $a$, $P_{2}$ to be the part of $P$ from $b^{\prime}$ to $s_{j}$, and $P_{3}$ to be the part of $P_{C}$ from $s_{j}$ to the first intersection with $P_{1}$. By a similar argument as in Case 1, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is a $(C,B;A)$-$\delta$-path.
Case 3: $a\in Q_{j}$ for some $j\neq i$, $c^{\prime}\in Q_{\ell}$ for some $\ell\neq i$.
Because of the clique on $X$, $S_{i}\in P_{A}\cap P_{B}\cap P_{C}$, so $i\in[k]$ and $P_{A}=P_{B}=P_{C}$, which also implies $\ell\in[k]$. If the order of these player-edges on $P_{B}$ is $B,A,C$, then we define $P_{1},P_{2},P_{3}$ as in Case 1. The argument is almost the same as Case 1, except that $P_{3}$ intersects $P_{1}$ because $S_{\ell}\in P_{1}\cap P_{B}$. As in Case 1, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is an $(A,B;C)$-$\delta$-path.
If the order of the player-edges in $P_{B}$ is $B,C,A$, then we define $P_{1}$ as the part of $P$ from $b$ to $a$, $P_{2}$ as the part of $P$ from $b^{\prime}$ to $s_{\ell}$, and $P_{3}$ as the part of $P_{B}$ from $s_{\ell}$ to the first intersection with $P_{1}$. $P_{3}$ intersects $P_{1}$ because $S_{j}\in P_{1}\cap P_{B}$. As in Case 2, $P_{1}\cup P_{2}\cup P_{3}\cup\{B\}$ is a $(C,B;A)$-$\delta$-path. ∎
Proposition 11.
Let $A,B,C\in\mathcal{K}$ and consider an $(A,B;C)$-$\delta$-path, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap B|=2$. Then $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$.
Proof.
First, we consider the case $|V(M)\cap C|=1$. Suppose for contradiction that $A$ and $B$ are on different alternating paths in $M\Delta M_{0}$: $P_{A}$ with other endpoint $x_{A}$, and $P_{B}$ with endpoints $x_{B}$ and $y_{B}$. Let $c^{\prime}$ be the tip of the $\delta$-path. We define $G^{\prime}$ by adding a complete graph on $x_{A},x_{B},y_{B},c^{\prime}$ and removing $A$. Then there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$, so we can consider the GE-decomposition (otherwise we get an alternating cycle through $A$ or a weakly blocking coalition).
The vertices $x_{A}$ and $c^{\prime}$ are in the same $Q_{i}$, and $i\in[k]$ because the $\delta$-path contains an alternating $a-c^{\prime}$ path and an alternating $a^{\prime}-c^{\prime}$ path. The edge $S_{i}$ is on both the path $P_{A}$ and the path of the $\delta$-path. We also know that $S_{i}$ is not in $P_{B}$, as the latter is disjoint from $P_{A}$. Since $B$ is on the cycle of the $\delta$-path, it is either in $Q_{j}$ for some $j\neq i$, or it is $S_{j}$ for some $j\neq i$. In both cases, either the $b-x_{B}$ or the $b^{\prime}-y_{B}$ subpath of $P_{B}$ has the property that it starts in $Q_{j}$ and does not contain $S_{j}$; we may assume by symmetry that it is the $b-x_{B}$ subpath. However, $x_{B}\in Q_{i}$ because of the edge $x_{A}x_{B}\in G^{\prime}$, so this subpath must contain $S_{i}$, a contradiction.
Next, consider the case $|V(M)\cap C|=2$. By Lemma 9, there is an $(A,B;C^{\prime})$-$\delta$-path such that $|V(M)\cap C^{\prime}|=1$. Using the first case for this $\delta$-path, we get that $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$. ∎
We define a subset $\mathcal{L}\subseteq\mathcal{K}$ the following way. A player $B$ is in $\mathcal{L}$, if for any $M_{0}$-alternating $A-B-C$ path for any $A\in\mathcal{K}$ and $C\in\mathcal{K}$, there exists an $(A,B;C)$-$\delta$-path or a $(C,B;A)$-$\delta$-path. Note that $\mathcal{K}^{*}\subseteq\mathcal{L}$ by Lemma 10.
For $A,B\in\mathcal{L}$, we say that $(A,B)$ is a pair if there exists an $(A,B;C)$-$\delta$-path for some $C\in\mathcal{K}$. If $(A,B)$ is a pair and $M\in\mathcal{SC}$ is a matching such that $|V(M)\cap A|=1$, then by Lemma 9 there is an $(A,B;C^{\prime})$-$\delta$-path such that $|V(M)\cap B|=2$ and $|V(M)\cap C^{\prime}|=1$. By Proposition 11, $A$ and $B$ are on the same alternating path in $M\Delta M_{0}$.
Lemma 12.
Let $A\in\mathcal{K}^{*}$, $B,C\in\mathcal{L}$ be such that $(A,B)$ and $(A,C)$ are pairs. Then $(B,C)$ is also a pair.
Proof.
Suppose for contradiction that $(B,C)$ is not a pair. Let $M\in\mathcal{SC}$ be a matching such that $|V(M)\cap A|=1$, and let $P$ be the path in $M\Delta M_{0}$ that contains $A$. By Proposition 11, $B$ and $C$ are on $P$ and $|V(M)\cap B|=|V(M)\cap C|=2$. We may assume that $P$ is an $A-B-C-D$ path, where $D\in\mathcal{K}^{*}$ and $|V(M)\cap D|=1$. Let the order of the vertices of these players along the path be $a,a^{\prime},b,b^{\prime},c,c^{\prime},d,d^{\prime}$.
Since $(A,C)$ is a pair, there is an $(A,C;E)$-$\delta$-path $H$ for some $E\in\mathcal{K}^{*}$ by Lemma 9 such that $|V(M)\cap E|=1$. Furthermore, since $C\in\mathcal{L}$ and $(B,C)$ is not a pair, we obtain that there exists a $(D,C;B)$-$\delta$-path, by considering the definition of $\mathcal{L}$ for the $B-C-D$ subpath of $P$. Using Lemma 9, we can get a $(D,C;A)$-$\delta$-path $H^{\prime}$, since the two ends of the path of $M\Delta M_{0}$ containing $B$ are $A$ and $D$, so $A$ must be the tip of the $\delta$-path given by Lemma 9.
We now consider the graph $G^{\prime}$ obtained from $E\cup M_{0}$ by adding $d^{\prime}e^{\prime}$ and removing $A$, where $e^{\prime}$ is the tip of $H$. Since there is no $M_{0}$-alternating $a-a^{\prime}$ path in $G^{\prime}$, we can consider the GE-decomposition. The vertices $d^{\prime}$ and $e^{\prime}$ are in the same $Q_{i}$, since there are $M_{0}$-alternating paths to them from $a$ or $a^{\prime}$, and $d^{\prime}e^{\prime}$ is an edge in $G^{\prime}$. We also know that $i\in[k]$, because $H$ contains an alternating $a-e^{\prime}$ path and an alternating $a^{\prime}-e^{\prime}$ path.
In $H^{\prime}$, the $D-A$ path containing $C$ and the $D-A$ path disjoint from $C$ both contain $S_{i}$. This means that $S_{i}$ is on the path of $H^{\prime}$, and $C$ is in $Q_{i}$. However, in $H$, both the $A-E$ path containing $C$ and the $A-E$ path disjoint from $C$ contain $S_{i}$. Thus, $S_{i}$ is on the path of $H$, and therefore $C$ is outside of $Q_{i}$, a contradiction. ∎
We are now ready to describe the structural properties that enable us to give a polynomial-time algorithm. Define $\mathcal{L}^{*}\subseteq\mathcal{L}$ the following way: a player $A\in\mathcal{L}$ is in $\mathcal{L}^{*}$ if for every $B,C\in\mathcal{L}$ such that $(A,B)$ and $(A,C)$ are pairs, $(B,C)$ is also a pair. It follows from Lemma 12 that $\mathcal{K}^{*}\subseteq\mathcal{L}^{*}$.
Since it can be decided in polynomial time whether an $(A,B;C)$-$\delta$-path exists for given $A,B,C\in\mathcal{K}$ (see Lemma 14), membership in $\mathcal{L}$ and $\mathcal{L}^{*}$ can be decided in polynomial time.
We introduce an auxiliary graph $G^{*}$ whose vertex set is $\mathcal{L}^{*}$ and $(A,B)$ is an edge if and only if it is a pair. This graph can be constructed in polynomial time, and, by the definition of $\mathcal{L}^{*}$, every component of $G^{*}$ is a clique. Let $\mathcal{K}_{0}$ denote the players $A\in\mathcal{K}$ from which there is no $M_{0}$-alternating path to any other player in $\mathcal{K}$. Obviously, members of $\mathcal{K}_{0}$ are isolated vertices in $G^{*}$ (but there may be other isolated vertices).
Let $\mathcal{Q}$ denote the set of components of $G^{*}$. For a component $Q\in\mathcal{Q}$, let $V_{Q}\subseteq V$ denote the set of vertices of all players in $Q$. The existence of a polynomial-time algorithm for finding a member of $\mathcal{SC}$ or determining that $\mathcal{SC}=\emptyset$ follows from the following theorem.
Theorem 13.
A matching $M$ of $G$ is in the strong core if and only if it satisfies the following conditions:
  • $|V(M)\cap A|=2$ for every $A\notin\mathcal{L}^{*}$,
  • $|V(M)\cap V_{Q}|\geq|V_{Q}|-1$ for every $Q\in\mathcal{Q}$, except when $Q=\{A\}$ and $A\in\mathcal{K}_{0}$.
Proof.
To see that the conditions are necessary, observe that $|V(M)\cap A|=2$ must hold for every $A\notin\mathcal{K}^{*}$. Also, if $(A,B)$ is a pair, then $|V(M)\cap(A\cup B)|\geq 3$ must hold by Lemma 9. Lastly, if $A\notin\mathcal{K}_{0}$, then $|V(M)\cap A|\geq 1$ must hold by Proposition 8.
To see sufficiency, we use the alternating path characterization in Lemma 5 for membership in $\mathcal{SC}$. Suppose $M$ satisfies the above two conditions. Then, the properties of $M$ imply that if a player-edge is in an $M_{0}$-alternating cycle, then both of its vertices are covered by $M$. Let $P$ be an $M_{0}$-alternating path from $A\in\mathcal{K}$ to $B\in\mathcal{K}$, where none of the vertices of $A$ are covered by $M$. Then $A\in\mathcal{K}_{0}$, which contradicts that there is an $M_{0}$-alternating path to some $B\in\mathcal{K}$ from $A$. Finally, suppose that $P$ is an $A-B-C$ alternating path, where $|V(M)\cap B|=1$. If $A\notin\mathcal{L}^{*}$, $B\notin\mathcal{L}^{*}$ or $C\notin\mathcal{L}^{*}$, then one of $A,B,C$ is covered twice by $M$, and we are done. Otherwise, since $A,C\in\mathcal{L}^{*}\subseteq\mathcal{K}$ and $B\in\mathcal{L}^{*}\subseteq\mathcal{L}$, the definition of $\mathcal{L}$ implies that $(A,B)$ is a pair or $(B,C)$ is a pair. In the former case, $|V(M)\cap A|=2$, while the latter case implies $|V(M)\cap C|=2$ by the second condition on $M$. Thus, the conditions of Lemma 5 are satisfied and therefore $M$ is in the strong core. ∎
Finding a matching with the properties in Theorem 13 can be done in polynomial time, since it amounts to finding a common base of a matching matroid and a generalized partition matroid. It remains to show that membership in $\mathcal{L}$ and $\mathcal{L}^{*}$ can be decided in polynomial time, which follows from the next lemma.
Lemma 14.
For given $A,B,C\in\mathcal{K}$, it can be decided in polynomial time if an $(A,B;C)$-$\delta$-path exists.
Proof.
We consider the graph $G^{\prime}$ obtained from $E\cup M_{0}$ by removing $A$. Since $A\in\mathcal{K}$, there is no $M_{0}$-alternating path between $a$ and $a^{\prime}$, so we can consider the GE decomposition. We may assume that $B$ has a vertex $b$ in some $Q_{i}$ and $C$ has a vertex $c$ in $Q_{j}$ for some $j\neq i$, since otherwise no $(A,B;C)$-$\delta$-path can exist. Since an $(A,B;C)$-$\delta$-path contains an alternating $a-c$ path and an alternating $a^{\prime}-c$-path, we may also assume $j\in[k]$.
Case 1: $i\in\{k+1,k+2\}$.
We may assume $i=k+1$ by symmetry. First, we find an $M_{0}$-alternating path $P_{1}$ from $a^{\prime}$ to $C$ (if no such path exists, then there is no $(A,B;C)$-$\delta$-path). Note that $P_{1}$ is disjoint from $Q_{i}$, and contains $S_{j}$.
Next, we check whether there is an $M_{0}$-alternating $a-s_{j}$ path that passes through $B$. To do this, we consider the graph $G^{\prime\prime}$ obtained from $G^{\prime}$ by adding $as_{j}$ and removing $B$ and $S_{j}$, and we check if there is an $(M_{0}\cup\{as_{j}\})$-alternating path from $b$ to $b^{\prime}$ in $G^{\prime\prime}$. Since no $M_{0}$-alternating cycle contains $B$, such a path must contain $as_{j}$, and therefore it gives an $M_{0}$-alternating path $P_{2}$ from $a$ to $s_{j}$ that traverses $B$. If $P_{2}$ does not exist, then there is no $(A,B;C)$-$\delta$-path.
Observe that $P_{2}$ must intersect $P_{1}$ because both contain $s_{j}$. Let $P^{\prime}_{2}$ be the subpath of $P_{2}$ up to the first intersection with $P_{1}$. Since $P_{1}$ is disjoint from $Q_{i}$, $P^{\prime}_{2}$ passes through $B$. Therefore, $P_{1}\cup P^{\prime}_{2}\cup\{A\}$ is an $(A,B;C)$-$\delta$-path.
Case 2: $i\in[k]$.
We check whether there are two vertex-disjoint $M_{0}$-alternating paths between $\{a,a^{\prime}\}$ and $\{s_{i},s_{j}\}$. We can do this by defining $M^{\prime}_{0}$ as $M_{0}\cup\{s_{i}s_{j}\}\setminus\{S_{i},S_{j}\}$, and checking whether there is an $M^{\prime}_{0}$-alternating path between $a$ and $a^{\prime}$. If no such path exists, then there is no $(A,B;C)$-$\delta$-path. Otherwise, by removing $s_{i}s_{j}$, we get a path $P_{1}$ from $a$ or $a^{\prime}$ to $s_{i}$, and a path $P_{2}$ from the other vertex of $A$ to $s_{j}$.
Next, we check whether there is an $M_{0}$-alternating $s_{i}-s_{j}$ path that starts with $S_{i}$ and passes through $B$. If $B=S_{i}$, then this simply means checking whether there is an alternating $s_{i}-s_{j}$ path that starts with $S_{i}$. Otherwise, we consider the graph $G^{\prime\prime}$ obtained from $G^{\prime}$ by adding $s^{\prime}_{i}s_{j}$ and removing $B,S_{i},S_{j}$, and checking whether there is an $(M_{0}\cup\{s^{\prime}_{i}s_{j}\})$-alternating path from $b$ to $b^{\prime}$ in $G^{\prime\prime}$. Since no $M_{0}$-alternating cycle contains $B$, such a path must contain $s^{\prime}_{i}s_{j}$, and therefore it gives an $M_{0}$-alternating path $P_{3}$ from $s_{i}$ to $s_{j}$ that starts with $S_{i}$ and traverses $B$. If $P_{3}$ does not exist, then there is no $(A,B;C)$-$\delta$-path.
We assume that $P_{3}$ exists. Observe that the only common vertex of $P_{1}$ and $P_{3}$ is $s_{i}$, otherwise $B$ would be in an $M_{0}$-alternating cycle. Also, $P_{3}$ must intersect $P_{2}$ because both end at $s_{j}$. Let $P_{3}^{\prime}$ be the subpath of $P_{3}$ up to the first intersection with $P_{2}$.
Finally, we find an $M_{0}$-alternating path $P_{4}$ from $s_{j}$ to $C$ that starts with $S_{j}$. If no such path exists, then there is no $(A,B;C)$-$\delta$-path. The path $P_{4}$ cannot leave $Q_{j}$, so it intersects $P_{1}\cup P_{2}\cup P^{\prime}_{3}$ only in $s_{j}$. Therefore, $P_{1}\cup P_{2}\cup P^{\prime}_{3}\cup P_{4}\cup\{A\}$ is an $(A,B;C)$-$\delta$-path. ∎
2.4 Examples
We present two examples in Figure 3 in order to explain the notions in the proof. The solid edges are in $M_{0}$, while the dashed edges are in $E$. In both examples, $\mathcal{K}=M_{0}$ because there is no $M_{0}$-alternating cycle, and $\mathcal{K}_{0}=\emptyset$. Also, $B\notin\mathcal{L}$ in both examples, since there is an $A-B-E$ alternating path, but there is neither an $(A,B;E)$-$\delta$-path nor an $(E,B;A)$-$\delta$-path. It can be checked that the remaining players are in $\mathcal{L}$ in both examples.
Figure 3: Solid edges are in $M_{0}$, dashed egdes are in $E$. The example on the right is obtained from the example on the left by adding $ab^{\prime}$. The strong core is empty on the left, while bold edges on the right form a matching in the strong core.
In the left example, the pairs are $(A,D),(A,E),(C,D),(C,E),(D,E)$, which means that =$\mathcal{L}^{*}=\{A,C,F\}$, and each are singletons in $G^{*}$. It can be checked that there is no matching that covers each of $A,C,F$ at least once and the remaining players twice, so $\mathcal{SC}=\emptyset$.
In the right example, $(A,C)$ is also a pair, so =$\mathcal{L}^{*}=\{A,C,D,E,F\}$, where $F$ is a singleton in $G^{*}$, while $A,C,D,E$ form a clique. By Theorem 13, $M\in\mathcal{SC}$ if and only if it covers both vertices of $B$, and it leaves at most one vertex of $F$ and at most one vertex of $A\cup C\cup D\cup E$ exposed. One such matching $M$ is indicated in bold; the dotted path is the single alternating path in $M\Delta M_{0}$.
3 Constant number of players
In this section, we show that if the number of players is bounded by a constant, then both the existence questions of a strong (resp. weak) core solution and the verification versions are solvable in polynomial time. This is relevant for example in the IKEP application, where players correspond to countries, which obviously implies a fairly small upper bound on the number of players. We shall see that the tractability of this case is a direct consequence of Lemma 1.
Recall that the instance is given by a graph $G=(V,E)$ and a partition $V=V_{1}\cup V_{2}\cup\dots\cup V_{m}$, where $m$ is the number of players. First, consider the verification problems. Lemma 1 immediately implies that if $m$ is constant, then it can be decided in polynomial time if a given matching $M$ is in the weak core. Indeed, the number of possible coalitions is polynomial, and for a given coalition $\{i_{1},i_{2},\dots,i_{k}\}$, we can decide, using Lemma 1, if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|+1$ for every $j\in[k]$.
Membership in the strong core can be decided similarly, but for every coalition $\{i_{1},i_{2},\dots,i_{k}\}$ and every $\ell\in[k]$, we check if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|$ for every $j\in[k]\setminus{\ell}$ and $|V(M^{\prime})\cap V_{i_{\ell}}|\geq|V(M)\cap V_{i_{\ell}}|+1$.
Corollary 15.
If the number of players is bounded by a constant $m$, then it can be decided in polynomial time whether a matching $M$ is in the weak (resp. strong) core.
Deciding non-emptiness of the weak and strong cores is somewhat more difficult, but still polynomial-time solvable. The crucial observation is that the membership of a matching $M$ in these cores depends only on the values $|V(M)\cap V_{i}|$ ($i\in[m]$). Using Lemma 1, we can find in polynomial time all component-wise maximal vectors $x\in\mathbb{Z}_{+}^{m}$ for which a matching $M$ with $|V(M)\cap V_{i}|\geq x_{i}$ for every $i\in[m]$ exists. Indeed, $x_{i}\leq|V_{i}|$ ($i\in[m]$) can be assumed, so we can check all possible vectors since $m$ is constant. Let $\mathcal{X}$ denote the set of such vectors. For each $x\in\mathcal{X}$, we can check in polynomial time, by using Lemma 1 for every coalition, whether any matching $M$ with $|V(M)\cap V_{i}|\geq x_{i}$ ($i\in[m]$) is in the strong or weak core. If the answer is negative for every $x\in\mathcal{X}$, then the core is empty, otherwise it is non-empty.
Corollary 16.
If the number of players is bounded by a constant $m$, then the emptiness of the weak (resp. strong) core can be decided in polynomial time. Furthermore, if the weak (resp. strong) core is nonempty, then a weak (resp. strong) core solution can also be found in polynomial time.
4 Hardness results for partition classes of constant size
In contrast to the case where the number of players is bounded by a constant, bounding the sizes of the partition classes by a constant does not lead to polynomial-time solvable problems. In this section, we show hardness results of this type. We have seen in Section 2 that our problems are tractable when all partition classes have size at most 2. Here we show that having partition classes of size 3 leads to coNP-hardness for most problems.
4.1 Weak core
In this section we focus on the weak core. Our first result is the hardness of verifying if a matching belongs to the weak core.
Theorem 17.
It is coNP-complete to decide if a matching $M$ is in the weak core, even if each partition class has size at most 3 and the graph is bipartite.
Proof.
We reduce from the NP-complete problem x3c (exact cover by 3-sets). An instance of x3c consists of $3n$ elements $\mathcal{X}=\{1,2,\dots,3n\}$ and a family of 3-sets $\mathcal{S}=\{S_{1},S_{2},\dots,S_{m}\}$ and the question is whether there exists $\mathcal{S^{\prime}}\subseteq\mathcal{S}$ such that the sets in $\mathcal{S^{\prime}}$ cover each element exactly once.
Let $I$ be an instance of x3c. We create a bipartite instance $I^{\prime}$ of the NTU international matching game as follows. First, we describe the vertex set $V$ of the graph.
  • For each element $i\in[3n]$ we create an element vertex $a_{i}$.
  • We also create connector vertices $b_{1,2}^{1},b_{1,2}^{2},b_{2,3}^{2},b_{2,3}^{3}\dots,b_{3n-1,3n}^{3n-1},b_{3n-1,3n}^{3n}$.
  • For each set $S_{j}\in\mathcal{S}$, we create vertices $c_{j}^{1},c_{j}^{2},c_{j}^{3}$ and vertices $d_{j}^{1},d_{j}^{2}$.
We now describe the edges of the graph.
  • For each $j\in[m]$, $l\in[2]$ we have an edge $c_{j}^{l}d_{j}^{l}$.
  • For each $i\in[3n-1]$, we have edges $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$.
  • For each set $S_{j}=\{j_{1},j_{2},j_{3}\}$ with $j_{1}<j_{2}<j_{3}$ we have edges $c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j_{3}}$.
The graph is bipartite, with bipartition $U=\{c_{1}^{1},c_{1}^{2},c_{1}^{3},\dots,c_{m}^{1},c_{m}^{2},c_{m}^{2}\}\cup\{b_{i,i+1}^{i}\mid i\in[3n-1]\}$ and $W=V\setminus U$.
Finally, we describe the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2}\}$. For each $i\in\{2,\dots,3n-1\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$. In addition, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, and $A_{3n}=\{b_{3n-1,3n}^{3n},a_{3n}\}$. Clearly, all players have size at most 3.
We create a matching $M$ as follows. We add $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[2]$ and $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i\in[3n-1]$.
Claim 18.
If there is an exact 3-cover in $I$, then $M$ is not in the weak core.
Proof.
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $\{C_{l_{1}},\dots,C_{l_{n}},A_{1},\dots,A_{3n}\}$ form a strongly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $(b_{i,i+1}^{i},b_{i,i+1}^{i+1})$ for $i=1,\dots,3n-1$. Hence, there is a matching $M^{\prime}$ among these players that matches every single one of their vertices. As each of these players had a vertex exposed in $M$ ($c_{j}^{3}$ for $C_{j}$, $a_{i}$ for $A_{i}$), it follows that $M$ is not in the weak core. ∎
Claim 19.
If $M$ is not in the weak core, then there is an exact 3-cover in $I$.
Proof.
Suppose that $M$ is not in the weak core. This means that there is a strongly blocking coalition $\mathcal{P}$ to $M$ with a matching $M^{\prime}$. We can suppose that there is a player $A_{i}\in\mathcal{P}$. Otherwise, there could be only $C_{j}$ players in $\mathcal{P}$ (the $D_{j}$ players have all vertices covered by $M$), and they induce no edges between them, contradiction.
Observe that players $A_{1},\dots,A_{3n}$ are all in $\mathcal{P}$, because for $A_{i}$ to strictly improve, $b_{i-1,i}^{i}$ and $b_{i,i+1}^{i}$ (if they exists) must be matched, so $b_{i-1,i}^{i-1}b_{i-1,i}^{i},b_{i,i+1}^{i}b_{i,i+1}^{i+1}\in M^{\prime}$. This implies that $A_{i-1},A_{i+1}\in\mathcal{P}$ and similarly we get that $A_{1},\dots,A_{3n}$ are all in $\mathcal{P}$.
Finally, as each player must strictly improve, every $a_{i}$ vertex is matched in $M^{\prime}$ and so they are matched to $C_{j}$ players. Whenever a $C_{j}$ player participates in the blocking by matching one of its vertices to an $a_{i}$ vertex, $\{c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j_{3}}\}\subseteq M^{\prime}$, because $D_{j}\notin\mathcal{P}$ and $C_{j}$ must strictly improve. Combining this with that fact that each $a_{i}$ is matched, we obtain that these players must correspond to an exact 3-cover. ∎
The statement of the theorem follows. ∎
Intuitively, it would be tempting to conjecture that the weak core is always nonempty in any NTU partitioned matching game. However, the following example (given by Zsuzsanna Jankó) shows that this is not the case.
Example 20.
Let $A=\{a_{1},a_{2},\dots,a_{7}\}$, $B=\{b_{1},\dots,b_{7}\}$, and $C=\{c_{1},\dots,c_{7}\}$ be three players each having seven vertices. The graph consists of five disjoint complete graphs, three $K_{5}$’s: $\{a_{1},c_{2},c_{3},b_{4},b_{5}\}$, $\{b_{1},a_{2},a_{3},c_{4},c_{5}\}$ and $\{c_{1},b_{2},b_{3},a_{4},a_{5}\}$, and two $K_{3}$’s: $\{a_{6},b_{6},c_{6}\}$ and $\{a_{7},b_{7},c_{7}\}$.
The weak core of this instance is empty. Indeed, if there were a matching in the weak core, then there would also be a maximum-size matching in it. Let $M$ be any maximum-size matching, covering 16 vertices. We may assume that every player has at least 4 vertices covered, since otherwise a single player would block $M$. The two players who have the least number of vertices covered have together at most $\frac{2}{3}\times 16<11$ vertices covered, and therefore, they have at most 10. If both of them have 5 covered in $M$, then they form a blocking coalition because there is a matching where both of them have 6 vertices covered. If one of them has 4 vertices covered while the other 6, then they form a blocking coalition because there is a matching where the first player has 5 vertices covered while the second has 7. Thus, every matching has a strongly blocking coalition.
We use this example in our gadgets to prove that deciding the emptiness of the weak core is hard.
Theorem 21.
It is NP-hard to decide whether a matching in the weak core exists, even if each partition class has size at most 7.
Before proving the theorem, we introduce the notion of special edges, which are actually not edges but gadgets that simulate edges with some special property. We say that there is a special edge between two vertices $u$ and $v$ if there is a gadget described below between $u$ and $v$. Let $H$ be a copy of the instance in Example 20. We have a player $S$ with four vertices $s_{1},s_{2},s_{3}$ and $s_{4}$. There is an edge between $s_{2}$ and $a_{1}\in H$. We also have a player $T$ with two vertices, $t_{1}$ and $t_{2}$ and edges $s_{1}t_{1}$, $s_{4}t_{2}$, $us_{1}$, $s_{4}v$ and $s_{2}s_{3}$. See Figure 4 for an illustration.
Figure 4: A special edge between $u$ and $v$.
Lemma 22.
Suppose an instance of the NTU partitioned matching game contains a special edge between $u$ and $v$. This special edge cannot belong to a matching in the weak core, but it can belong to a strongly blocking coalition in the following sense:
  1. i)
    If there is a matching $M$ in the weak core, then $s_{2}a_{1}$, $s_{1}t_{1}$ and $s_{4}t_{2}$ has to belong to $M$. This means that $us_{1}$ and $s_{4}v$ cannot belong to $M$.
  2. ii)
    Suppose the edges $us_{1}$ and $s_{4}v$ do not belong to a matching $M$. If by replacing the special edge between $u$ and $v$ with an edge $uv$, the players containing $u$ and $v$ would belong to a strongly blocking coalition w.r.t. $M$ restricted to the graph we get if we delete the special edge, then $M$ does not belong to the weak core.
Proof.
i) We have seen that if there are no additional edges, then $H$ cannot admit a matching in the weak core. Therefore if $s_{2}a_{1}$ does not belong to $M$, then the whole instance could not admit a matching in the weak core.
Suppose $s_{1}t_{1}$ or $s_{4}t_{2}$ does not belong to $M$, so $t_{1}$ or $t_{2}$ is unmatched in $M$. We know that $s_{2}a_{1}$ belongs to $M$, therefore $s_{3}$ is unmatched in $M$. But this means that the players $S$ and $T$ with the matching $M^{\prime}=\{s_{1}t_{1}$, $s_{2}s_{3}$, $s_{4}t_{2}\}$ form a blocking coalition, because $M^{\prime}$ covers every vertex of $S\cup T$.
ii) We have already seen that if the edges $s_{1}t_{1}$, $s_{2}a_{1}$ and $s_{4}t_{2}$ do not belong to $M$, then $M$ is not in the weak core. Suppose these edges belong to $M$, so $s_{3}$ is unmatched in $M$. Let $M^{\prime}$ be the blocking matching. Let $M^{\prime\prime}$ be the matching we get if we leave out the edge $uv$ from $M^{\prime}$ and add the edges $us_{1}$, $s_{4}v$ and $s_{2}s_{3}$. The players that belonged to the strongly blocking coalition would also be better off by $M^{\prime\prime}$, so together with $S$, they form a strongly blocking coalition: $S$ prefers $M^{\prime\prime}$ because all vertices of $S$ are covered by $M^{\prime\prime}$. ∎
Proof of Theorem 21.
We reduce from 3-SAT. Given an instance $I$ of 3-SAT, we construct an instance $I^{\prime}$ of the NTU partitioned matching game. For every variable of $I$, we construct a variable gadget, and for every clause we construct a clause gadget. After that, we describe the interconnecting edges between the clause gadgets and the variable gadgets.
Variable gadget. For a variable $x_{i}$, we define four sets of vertices, $X_{i}$, $\bar{X_{i}}$, $Y_{i}$ and $\bar{Y_{i}}$. $X_{i}$ contains a vertex for every occurrence of the variable $x_{i}$ in unnegated form, and the vertices of $X_{i}$ belong to one player. $\bar{X_{i}}$ contains a vertex for every occurrence of the variable $x_{i}$ in negated form, and the vertices of $\bar{X_{i}}$ belong to one player. For every vertex in $X_{i}$, there is a separate vertex in $Y_{i}$ so that there is an edge between them, and each vertex of $Y_{i}$ belongs to a separate player of size one. Similarly, for every vertex in $\bar{X_{i}}$, there is a separate vertex in $\bar{Y_{i}}$ so that there is an edge between them, and each vertex of $\bar{Y_{i}}$ belongs to a separate player of size one. Every vertex of $Y_{i}$ is connected to every vertex of $\bar{Y_{i}}$ by a special edge. See Figure 5.
Figure 5: The variable gadget. The dotted lines represent special edges.
Clause gadget. For each clause $c_{j}$, we associate a copy of the instance in Example 20. We use the same notation for the vertices as in the example, but every vertex gets an upper index $j$. Let $a_{1}^{j}$, $b_{1}^{j}$ and $c_{1}^{j}$ correspond to the three literals in clause $c_{j}$.
Interconnecting edges Now we describe the edges between the clause gadgets and the variable gadgets. To every literal that appears in a clause, there is a corresponding vertex in the clause gadget and a corresponding vertex in the variable gadget. We connect these two vertices with an edge.
Claim 23.
For any matching $M$ in the weak core of $I^{\prime}$, there is a truth assignment satisfying $I$.
Proof.
If $M$ is in the weak core, then $Y_{i}$ or $\bar{Y_{i}}$ has all its vertices covered by $M$. Indeed, if there were a vertex $u\in Y_{i}$ and a vertex $v\in\bar{Y_{i}}$ so that neither is covered by $M$, then the players $\{u\}$ and $\{v\}$ together with the special edge connecting them would form a strongly blocking coalition (see Lemma 22). Since the special edges cannot belong to $M$, every vertex of $X_{i}$ (or $\bar{X_{i}}$) is matched to a vertex in $Y_{i}$ (or $\bar{Y_{i}}$). In the first case we set the variable $x_{i}$ to be false, and in the second case we set it to be true. If both hold simultaneously, we arbitrarily set $x_{i}$ to be true or false, but it still holds that if $x_{i}$ is set to false, every vertex in $X_{i}$ is matched to a vertex in $Y_{i}$ in $M$.
A clause gadget is a copy of the instance in Example 20, therefore without the additional (interconnecting) edges, it cannot admit a matching in the weak core, thus $M$ has to contain an interconnecting edge leaving this clause gadget. The interconnecting edge corresponds to a literal that cannot be set to false, because then the vertex corresponding to it in $X_{i}$ or $\bar{X_{i}}$ for some $i$ would be matched to a vertex in $Y_{i}$ or $\bar{Y_{i}}$ in $M$. This means that in this truth assignment, every clause contains a literal set to true, thus it satisfies $I$. ∎
Claim 24.
For any truth assignment satisfying $I$, it is possible to construct a matching $M$ in the weak core of $I^{\prime}$.
Proof.
If a variable $x_{i}$ is set to true (false), let $M$ contain all the edges between $\bar{X_{i}}$ ($X_{i}$) and $\bar{Y_{i}}$ ($Y_{i}$), and all interconnecting edges incident to $X_{i}$ ($\bar{X_{i}}$), see Figure 6.
Figure 6: Red edges are in $M$. The dotted lines represent special edges.
In a special edge $uv$, let $s_{1}t_{1}$, $s_{2}a_{1}$, $s_{4}t_{2}$ and $c_{2}c_{3}$, $b_{4}b_{5}$, $a_{2}a_{3}$, $c_{4}c_{5}$, $b_{2}b_{3}$, $a_{4}a_{5}$, $b_{6}c_{6}$, $b_{7}c_{7}$ belong to $M$, see Figure 7.
Figure 7: The red edges belong to $M$.
In every clause gadget, there is at least one vertex corresponding to a literal that is matched via an interconnecting edge in $M$ since every clause contains a literal set to true. Suppose that $a_{1}^{j}$ is such a vertex in the clause gadget corresponding to $c_{j}$. Let the edges $c_{2}^{j}c_{3}^{j}$, $b_{4}^{j}b_{5}^{j}$, $a_{2}^{j}a_{3}^{j}$, $c_{4}^{j}c_{5}^{j}$, $b_{2}^{j}b_{3}^{j}$, $a_{4}^{j}a_{5}^{j}$, $b_{6}^{j}c_{6}^{j}$ and $b_{7}^{j}c_{7}^{j}$ belong to $M$.
Now we show for every player that it cannot belong to a strongly blocking coalition. If the variable $x_{i}$ is set to true, then all the vertices in $X_{i}$, $\bar{X_{i}}$ and $\bar{Y_{i}}$ are matched in $M$, therefore the players $X_{i}$, $\bar{X_{i}}$ and the singleton players of $\bar{Y_{i}}$ cannot belong to a strongly blocking coalition. Since a vertex (which is also a player) in $Y_{i}$ is only connected to one of these players, it also cannot belong to a blocking coalition. If $x_{i}$ is set to false, a similar argument shows that these players cannot belong to a strongly blocking coalition.
Next, we show that a player of a special edge cannot belong to a strongly blocking coalition. $T$ cannot, since it is fully matched in $M$. $S$ cannot, since it has only one vertex unmatched in $M$, and $s_{1}$ is only connected to a vertex in $Y_{i}$ or $\bar{Y_{i}}$ and a vertex in $T$, therefore it would be unmatched in any blocking coalition. It remains to check that some of the players $A$, $B$ and $C$ do not form a blocking coalition. This is because the maximum matchings in $A$, $B$, $C$, $A\cup B\cup C$ and $B\cup C$ cover at most as many vertices as $M$, and a maximum matching in $A\cup B$ (or $A\cup C$) covers 12 vertices while $M$ covers 11 vertices of $A\cup B$ (and $A\cup C$).
This last argument about the players of the example shows that the players of a clause gadget also cannot belong to a strongly blocking coalition, which completes the proof. ∎
This concludes the proof of Theorem 21.∎
4.2 Strong core
We proceed to prove analogous hardness results for the strong core.
Theorem 25.
It is coNP-complete to decide if a matching $M$ is in the strong core, even if each partition class has size at most 3. Furthermore, it is NP-hard to decide if the strong core is empty, even if each partition class has size at most 3. Both hardness results hold for bipartite graphs.
Proof.
We reduce from the NP-complete problem x3c as in the proof of Theorem 17; we use the same notation as in that proof. Let $I$ be an instance of x3c. We create an instance $I^{\prime}$ of the international matching game as follows. First we describe the vertex set $V$ of the graph.
  • For each element $i\in[3n]$ we create an element vertex $a_{i}$ and a friend vertex $x_{i}$, except that we do not create $x_{3n}$ (only $x_{1},\dots,x_{3n-1}$).
  • We also create connector vertices $b_{1,2}^{1},b_{1,2}^{2},b_{2,3}^{2},b_{2,3}^{3}\dots,b_{3n-2,3n-1}^{3n-2},b_{3n-2,3n-1}^{3n-1}$ and $y_{1,2}^{1},y_{1,2}^{2},y_{2,3}^{2},$ $y_{2,3}^{3},\dots,y_{3n-2,3n-1}^{3n-2},y_{3n-2,3n-1}^{3n-1}$.
  • For each set $S_{j}\in\mathcal{S}$, we create vertices $c_{j}^{1},c_{j}^{2},c_{j}^{3}$ and vertices $d_{j}^{1},d_{j}^{2},d_{j}^{3}$.
We now describe the edges of the graph.
  • For each $j\in[m]$, $l\in[3]$ we have an edge $c_{j}^{l}d_{j}^{l}$.
  • For each $i\in[3n-1]$ we have an edge $x_{i}a_{i}$ (note that there is no $x_{3n}$ vertex).
  • For each $i\in[3n-2]$, we have edges $b_{i,i+1}^{i}b_{i,i+1}^{i+1},y_{i,i+1}^{i}y_{i,i+1}^{i+1}$.
  • For each set $S_{j}=\{j_{1},j_{2},j_{3}\}$ with $j_{1}<j_{2}<j_{3}$ we have edges $c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j}^{3}$.
The graph is bipartite with bipratition $U=\{c_{1}^{1},c_{1}^{2},c_{1}^{3},\dots,c_{m}^{1},c_{m}^{2},c_{m}^{2},x_{1},\dots,x_{3n-1}\}\cup\{b_{i,i+1}^{i},y_{i,i+1}^{i}\mid i\in[3n-1]\}$ and $W=V\setminus U$.
Finally, we describe the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2},d_{j}^{3}\}$. For each $i\in\{2,\dots,3n-2\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$ and a player $X_{i}=\{y_{i-1,i}^{i},x_{i},y_{i,i+1}^{i}\}$ Finally, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, $X_{1}=\{x_{1},y_{1,2}^{1}\}$ and $A_{3n-1}=\{b_{3n-2,3n-1}^{3n-1},a_{3n-1},a_{3n}\}$, $X_{3n-1}=\{y_{3n-2,3n-1}^{3n-1},x_{3n-1}\}$. All players have size at most 3.
We start by showing that if there is a matching $M$ in the strong core of $I^{\prime}$, then $M$ is the matching that contains all edges of $E$, except the ones of the form $c_{j}^{l}a_{j_{l}}$. First, observe that $M$ must contain all edges $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[3]$. Suppose that there is an edge $c_{j}^{l}d_{j}^{l}$ not in $M$. Then, player $D_{j}$ does not have all its vertices covered in $M$. However, $C_{j}$ and $D_{j}$ can create a matching among themselves that matches each of their vertices, so $\{C_{j},D_{,}\}$ would be a weakly blocking coalition to $M$, a contradiction. This implies that none of the edges of the form $c_{j}^{l}a_{j_{l}}$ are in $M$, as the $c_{j}^{l}$ vertices are matched elsewhere. Finally, observe that all the remaining edges can be added simultaneously, hence $M$ must contain all of them, as any strong core matching is maximal. Note that this matching $M$ matches every vertex of $V$, except $a_{3n}$.
Therefore, by showing that this matching $M$ is in the strong core if and only if there is no exact 3-cover, we show the desired hardness of both problems.
Claim 26.
If there is an exact 3-cover in $I$, then $M$ is not in the strong core.
Proof.
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $C_{l_{1}},\dots,C_{l_{n}},A_{1},A_{2},\dots,A_{3n-1}$ form a weakly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i=1,\dots,3n-2$. Hence, there is a matching $M^{\prime}$ among these players that matches every single one of their vertices. As $A_{3n-1}$ had a vertex exposed in $M$, it follows that $M$ is not in the strong core. ∎
Claim 27.
If $M$ is not in the strong core, then there is an exact 3-cover in $I$.
Proof.
Suppose that $M$ is not in the strong core. This means that there is a weakly blocking coalition $\mathcal{P}$ to $M$ with a matching $M^{\prime}$ in the induced graph $G[\mathcal{P}]$. As $a_{3n}$ is the only vertex that is not covered by $M$, it follows that player $A_{3n-1}$ is in the blocking coalition and the vertex $a_{3n}$ is matched. As it can only be matched to a player $C_{j}$ such that $a_{3n}\in S_{j}$, it follows that there is a player $C_{j}$ in the coalition too. Furthermore, one of the vertices of $C_{j}$ is matched to $a_{3n}$, hence $D_{j}$ cannot have all its vertices covered and therefore cannot be in the coalition. As $C_{j}$ must weakly improve, all of its vertices must be covered by $M^{\prime}$. Hence, $c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{3n}\in M^{\prime}$ and thus $A_{j_{1}},A_{j_{2}}\in\mathcal{P}$.
Now, we make two observations. The first is that none of the players $X_{1},\dots,X_{3n-1}$ can participate in $\mathcal{P}$. Suppose for the contrary that $X_{i}\in\mathcal{P}$. Then, $M^{\prime}$ covers $y_{i-1,i}^{i},y_{i,i+1}^{i}$ (whenever such a vertex exists). Hence, $X_{i-1},X_{i+1}\in\mathcal{P}$ and by iterating this argument we get that $X_{i}\in\mathcal{P}$ for $i\in[3n-1]$. However, $a_{j_{1}}$ is not matched to $x_{j_{1}}$, hence $x_{j_{1}}$ is exposed in $M^{\prime}$, so player $X_{j_{1}}$ does not weakly improve, contradiction. The second observation is that players $A_{1},\dots,A_{3n-1}$ are all in $\mathcal{P}$. We have already seen that $A_{3n-1}\in\mathcal{P}$. Therefore, to weakly improve, $b_{3n-2,3n-1}^{3n-1}$ is matched, so $b_{3n-2,3n-1}^{3n-1}b_{3n-2,3n-1}^{3n-2}\in M^{\prime}$. This implies that $A_{3n-2}\in\mathcal{P}$ and by similar arguments we get that $A_{3n-3},\dots,A_{1}$ are in $\mathcal{P}$.
Finally, as each player must weakly improve, every $a_{i}$ vertex is matched in $M^{\prime}$ and by the first observation they are matched to $C_{j}$ players. Whenever a $C_{j}$ player participates in the blocking by matching one of its vertices to an $a_{i}$ vertex, its corresponding $D_{j}$ player cannot keep all its vertices covered, so it is not in the blocking coalition. This implies that whenever $C_{j}\in\mathcal{P}$, $\{c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j_{3}}\}\subseteq M^{\prime}$. Combining this with the fact that each $a_{i}$ is matched, we obtain that these players must correspond to an exact 3-cover. ∎
This completes the proof of the theorem. ∎
Acknowledgement
The authors would like to thank Péter Biró and Zsuzsanna Jankó for the fruitful discussions on the topic. Gergely Csáji is supported by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858 and by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation fund, financed under the KDP-2023 funding scheme (grant number C2258525). Tamás Király is supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant number ELTE TKP 2021-NKTA-62 and grant number K143858.